RPC weekly meeting

Europe/Zurich
CERN

CERN

Davide Boscherini (Universita e INFN, Bologna (IT)), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Mohamed Zaazoua (University of Science and Technology of China (CN))
Description

Weekly meeting about the RPC system activities

Zoom Meeting ID
61128581037
Host
Davide Boscherini
Useful links
Join via phone
Zoom URL
    • 09:15 09:25
      Communication to/from gas group 10m
      Speakers: Beatrice Mandelli (CERN), Enrico Pastori (INFN e Universita Roma Tor Vergata (IT)), Gianluca Rigoletti (CERN), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Kacper Kapusniak (CERN), Roberto Guida (CERN)
    • 09:25 09:35
      Detector and trigger reports (incl. BIS78) 10m
      Speakers: Alessandro Polini (Universita e INFN, Bologna (IT)), Claudio Luci (Sapienza Universita e INFN, Roma I (IT)), Enrico Pastori (INFN e Universita Roma Tor Vergata (IT)), Enrico Tusi (INFN e Universita Roma Tor Vergata (IT)), Eric Ballabene (University and INFN, Bologna (IT)), Dr Federico Lasagni Manghi (Universita e INFN, Bologna (IT)), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Mr Mohamed Zaazoua (University of Science and Technology of China (CN)), Sinem Simsek (Istinye University (TR))

      BMS.C.10.CO.Ly0 HV channel 

      • SPC Gas gap firstly masked without reducing the HV of the HV channel [e-log]
      • Gas gap driving high current also contributing to high rate of s17_t4 (1.2kHz instead of ~1Hz).
        BMS5C10.PI.Ly0.ETA Vth lowered from 0.7 to 0.3 [e-log].
      • Gas gap still driving high current, HV lowered to 8000V [e-log]
      • Gas gap still driving high current, HV lowered to 5000V and gap unmasked [e-log]
      • Cavern intervention: chamber disconnected, HV channel brought back to nominal [e-log]

       

      RPC BMS.A.16.CO.Ly1 HV channel

      • SPC Gas gap driving high current (60uA at 9056V).
        Already at 5000V, the SPC gas gap was driving high current (32uA).
        HV channel lowered and masked. [e-log]
      • Cavern intervention: chamber disconnected, HV channel brought back to nominal [e-log]

       

      RPC HV channels in S11/S12

      • 6 RPC HV channels in sectors 11 and 12 went OFF [e-log]
        The HV board went into Undervoltage state (with Vmon ~100V), lowering the HV for all 6 connected channels. 
        The board was responding to DCS but not setting the HV consistently.
        The board had been switched off, and the corresponding channels had been masked. 
      • Cavern intervention: HV board giving Undervoltage error has been replaced with a spare [e-log]

       

      Cavern intervention on 19/11 [e-log]

      • See the previous points.

       

      BOS5C08 Vpad iCalib set to 11.0 from 15.7 because s13_t4 not reachable [e-log, e-log

      • While ATLAS was running without beams in the machine, a TTC restart was performed, which did not resolve the problem.
        After that, the trigger was put on hold and a power cycle performed from DCS on the two RPC PADs BMS5C08 and BOS5C08, followed by a second TTC restart. The issue persisted. It therefore seems to be a hardware issue that cannot be fixed without access to the cavern. It might also be related to the toroid ramp-dow.

       

      BML4C07.CO Vee [e-log]

      • RPC was in NOT_READY due to an error of BML4C07.CO Vee.
        BML4C07.CO Vee at 1.8A instead of 2.2A -- masked.
        BML4C07.CO Vpd, that was already masked, gives OverCurrent.
    • 09:35 09:45
      Plan for YETS 10m
      Speakers: Ahmet Renklioglu (Istinye University (TR)), Alessandro Polini (Universita e INFN, Bologna (IT)), Claudio Luci (Sapienza Universita e INFN, Roma I (IT)), Dmitri Kharchenko (Joint Institute for Nuclear Research (RU)), Enrico Pastori (INFN e Universita Roma Tor Vergata (IT)), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Uladzimir Kruchonak (Joint Institute for Nuclear Research (RU))